# Pierce the Mists, Greet the Sky: Decipher Knowledge Overshadowing via Knowledge Circuit Analysis

### v2

@@ Abstract
Large Language Models (LLMs), despite their remarkable capabilities, are hampered by hallucinations. A particularly challenging variant, knowledge overshadowing, occurs when one piece of activated knowledge inadvertently masks another relevant piece, leading to erroneous outputs even with high-quality training data. Current understanding of overshadowing is largely confined to inference-time observations, lacking deep insights into its origins and internal mechanisms during model training. Therefore, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect knowledge overshadowing. By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the internal workings of attention heads, tracing how competing knowledge pathways contribute to the overshadowing phenomenon and its evolution throughout the training process. Extensive experiments demonstrate PhantomCircuit ’s effectiveness in identifying such instances, offering novel insights into this elusive hallucination and providing the research community with a new methodological lens for its potential mitigation.
@@
Haoming Huang1,3, [^1] Co-first authors with equal contribution., Yibo Yan1,2, [^1] Co-first authors with equal contribution., Jiahao Huo1, Xin Zou1, Xinfeng Li4, Kun Wang4, Xuming Hu1,2, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University
@@ 2.2 Knowledge Circuit Analysis
In the context of mechanistic interpretability Rai et al. (2024); Huo et al. (2024); Huang et al. (2024a), computations in Transformer-based language models are viewed as a connected directed acyclic graph encompassing components such as MLPs and attention layers Syed et al. (2023); Conmy et al. (2023); Huang et al. (2024b). A circuit refers to a sparse computational subgraph that significantly influences the model’s behavior on a specific task Olah et al. (2020); Elhage et al. (2021); Wang et al. (2022). Building on this, Yao et al. (2024) introduce the concept of knowledge circuits, hypothesizing that cooperation among components reveals implicit knowledge in LLMs. Further, Ou et al. (2025) explore how such circuits evolve during continual pre-training, providing insights into knowledge acquisition. To enable effective knowledge editing, CaKE Yao et al. (2024) proposes a Circuit-aware Knowledge Editing method that guides models to activate modified knowledge and form new reasoning pathways. In this paper, we analyze the phenomenon of knowledge overshadowing through the lens of knowledge circuits, contributing new perspectives to LLM hallucination detection.
@@ 3.1 Knowledge Overshadowing
Let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$. The core idea is that a strong learned pattern $X_{dom}\leftrightarrow X_{bg}$ can "over-generalize" where the model primarily associates $X_{dom}$ with $X_{bg}$.
Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$), it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
@@ 3.1.1 Knowledge Overshadowing Occurrence
When the input prompt is $P_{sub}$, but the model wrongly generates $Y_{dom}$, knowledge overshadowing occurs for this query-response instance, resulting in the $(P_{sub},Y_{dom})$.
@@ 3.1.2 Quantitative Indicators
Let $N_{dom}$ and $N_{sub}$ be the number of instances of $P_{dom}$ and $P_{sub}$ in the training dataset, respectively. The dataset $Z$ comprises a subset $Z_{dom}$ containing $N_{dom}$ instances of $P_{dom}$, and a subset $Z_{sub}$ containing $N_{sub}$ instances of $P_{sub}$.
@@ 3.1.3 Overshadowing Influence Factors
Knowledge Popularity (P) is the fundamental cause of the knowledge overshadowing phenomenon and serves as its primary influencing factor. P is defined as the ratio of the capacities of $Z_{dom}$ to $Z_{sub}$, thus $P=N_{dom}/N_{sub}.$
@@
In addition to the factors mentioned in Zhang et al. (2025a, 2024b), aiming to analyze the dynamic evolution of knowledge overshadowing during the training process, we extend our consideration to total number of tokens, the Dataset Size (D) in the training set. The average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch is defined as $\mathcal{LP}=loss(P_{sub})/total\;loss,$ which also relates to the overshadowing.
@@ 3.2.2 Knowledge Circuit Construction
We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs, $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, $E$ represents the information flow between these components. Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of given input prompt, particularly in knowledge overshadowing, is $\{X_{dom},X_{sub}\}$, the difference between $P_{dom}$ and $P_{sub}$. The adapted construction methods is similar to edge attribution patching Conmy et al. (2023), which involves:
Paired inputs. For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrasting mentioned below.
Activation difference calculation. After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activation of parent node when $P_{sub}$ as clean input and $P_{dom}$ as corrupt input.
Edge score calculation. The importance $S(e)$ of an edge $e=(v_{p},v_{c})$ is scored based on how patching $v_{p}$’s activation (using $\Delta A(v_{p})$) influences a metric $\mathcal{M}$, which assesses the model’s ability to correctly output $Y_{sub}$ rather than $Y_{dom}$ for $P_{sub}$ inputs. Using the Integrated Gradients, $S(e)$ is approximated as:
@@ 3.2.3 Circuit-based Analysis
Node Attention Analysis. We identify high attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
Circuit Structure Analysis. We then trace the information flow of these high attention heads by identifying their parent and child nodes to understand their structural role. Nodes consistently retained in circuits built with different thresholds $\tau$ are also analyzed as key components.
@@ 3.2.4 Circuit-based Overshadowing Recovery
where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$, This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
@@
Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, is masked. $Y_{dom}$ is identified as $y_{i}$ that has the highest average rank across all $P^{\prime}_{sub}$. For circuit construction, $X_{dom}$ within $P_{dom}$ is replaced by a generic placeholder like "something", as shown in Figure 5. Combining this streamlined approach enables broader application of our knowledge circuit-based overshadowing recovery. See more details in Appendix B.2
@@ 4.1 Experiment Setup
Figure 3: The co-evolution between $\mathcal{RO}$ $\mathcal{LP}$ in different M. Early high overall loss and low $\mathcal{LP}$ leads to intensive optimization of $P_{dom}$ and $\mathcal{RO}$ rises up to near 100%. As training progresses, subordinate knowledge loss proportion ($\mathcal{LP}$) rises, shifting optimization focus to $P_{sub}$ errors, initiating $\mathcal{RO}$’s recovery phase, validated across models.
@@ 4.1.1 Dataset
Finetuning Dataset. We construct a finetuning dataset to evaluate circuit-base overshadowing recovery method. Utilizing virtual knowledge, we preserve the natural language semantics while avoiding prior knowledge editing that could interfere with the natural occurrence of overshadowing. For this dataset, we set P = 5 and D = 1M. The Appendix D shows the dynamics analysis based on finetuning dataset.
@@ 4.1.4 Evaluation
To measure $\mathcal{RO}$, we randomly sample 500 $P_{dom}$ and 500 $P_{sub}$ prompts for evaluation after each training epoch. The $\mathcal{LP}$ is recorded within each epoch. The results are shown in Figures 2 and 3.
Figure 4: The circuit-based analysis of proposed PhantomCircuit. (a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch. (b) focuses on the onset and recovery phases and shows that the appear and disappear of high attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$. (C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers. The main structure of circuit (with 400 edges totally) and attention mode show that high attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank. Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
@@ 4.2 Main Result
As shown in Figures 2(a) and 2(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase. This is attributed to more prominent dominant patterns $X_{dom}\leftrightarrow X_{bg}$ being learned and generalized rapidly, even within the first epoch. The duration phase also decreases with higher P, because a larger P and a fixed D implies fewer groups of knowledge pairs, leading to less diversity of $P_{sub}$, allowing the model to learn all overshadowed $P_{sub}$ and recover from overshadowing more quickly.
Figures 2(b) and 2(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing. This is due to the stronger generalization capabilities of larger models, leading them to quickly learn and overgeneralize dominant patterns. However, larger models also show a shortened duration phase and a rapid decline in $\mathcal{RO}$ during recovery, suggesting enhanced capacity to differentiate $\{X_{dom},X_{sub}\}$, thus recovering faster despite earlier overshadowing.
In Figures 2(c) and 2(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner. This is because more data per epoch provides more iterations and exposure to the dominant pattern, accelerating its generalization. Conversely, the recovery phase is prolonged with larger datasets. The increased diversity of $P_{sub}$ in larger datasets requires more epochs for the model to learn all instances and recover.
Figure 5: Overshadowing recovery via optimized circuit. In the finetuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$. First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
@@
The occurrence of overshadowing in pretrained LLMs can be understood with dynamics analysis results above. Initially, large M and D promote the rapid generalization of $X_{dom}\leftrightarrow X_{bg}$ and overshadowing onset. The subsequent inadequate optimization of the $P_{sub}$ is exacerbated by the sheer scale and diversity of training data, which brings prolonged overshadowing effect and exceeded sharp recovery effect from large M (e.g., trillion tokens for training LlaMa-2-7B ).
@@
The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs dictates the extent of knowledge overshadowing. Figure 4 (a) shows the variation of circuit’s average attention scores on $\{X_{dom},X_{sub}\}$ throughout the training phase. The higher average attention alleviates overshadowing, while lower attention exacerbates it.
@@
Figure 4 (b) shows the circuit dynamics across the onset and recovery phases of knowledge overshadowing. These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit. The threshold for high attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset. Conversely, when knowledge overshadowing intensifies, a subset of these critical attention heads tends to disappear from the circuit.
@@
Furthermore, by focusing on the circuit’s internal mechanisms and structure within a specific epoch, as shown in Figure 4 (c), we leverage the circuit-based analysis of our proposed PhantomCircuit. First, according to the logits and ranks of $Y_{dom}$ and $Y_{sub}$ across layers, the 5th layer is a juncture where the rank of $Y_{sub}$ exceeds $Y_{dom}$. Subsequently, by focusing on the circuit structure, we identify the internal mechanisms driving this juncture. A high attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5. The a5.h1 also appears to inherit its high attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal. The attention map on the right confirms its high focus on $\{X_{dom},X_{sub}\}$.
@@
Knowledge circuit-guided optimization represents a promising strategy to mitigate the knowledge overshadowing effect. Figure 5 illustrates our findings. We first finetune the model on finetuning dataset. During the recovery phase, we randomly choose some $P_{sub}$. Feeding $P_{sub}$ to the original full model (M = 410M), the prediction is still $Y_{dom}$ instead of expected $Y_{sub}$, and shows a significant gap in probability as well.
@@
We then detect the $Y_{sub},Y_{dom}$ and $X_{sub}$ for chosen $P_{sub}$ automatically, replace the $X_{sub}$ of placeholder token "something". With these components, we obtain the $P_{dom}$ as corrupt input and $P_{sub}$ as clean input to construct a knowledge circuit. A golden section search algorithm is employed to determine the optimal number of edges for building $C_{opt}$. The optimized circuit structure map shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information. Retained high attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$. Some low attention heads also remain, implying that even in the circuit, processing background knowledge $X_{bg}$ and linking it to the distinctive elements $X_{sub}$ is crucial for correct inference. The performance of the optimized circuit $C_{opt}$ is then evaluated by feeding it the clean input $P_{sub}$, while $P_{dom}$ serves as the baseline for contrast. Finally, the circuit successfully produces $Y_{sub}$, demonstrating the elimination of the overshadowing effect.
Future work will enhance the circuit optimization metric $\mathcal{M}$ by incorporating $Y_{sub}$’s absolute logit alongside the logit difference with $Y_{dom}$ for more effective guidance. Developing a comprehensive evaluation framework for circuit-based recovery is also crucial. These steps will evolve PhantomCircuit into an integrated platform for efficient analysis and robust optimization of knowledge overshadowing.

### v4

@@ Abstract
Large Language Models (LLMs), despite their remarkable capabilities, are hampered by hallucinations. A particularly challenging variant, knowledge overshadowing, occurs when one piece of activated knowledge inadvertently masks another relevant piece, leading to erroneous outputs even with high-quality training data. Current understanding of overshadowing is largely confined to inference-time observations, lacking deep insights into its origins and internal mechanisms during model training. Therefore, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect knowledge overshadowing. By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the function of key components in the circuit and how the attention pattern dynamics contribute to the overshadowing phenomenon and its evolution throughout the training process. Extensive experiments demonstrate PhantomCircuit ’s effectiveness in identifying such instances, offering novel insights into this elusive hallucination and providing the research community with a new methodological lens for its potential mitigation. Our code can be found in https://github.com/halfmorepiece/PhantomCircuit.
@@
Haoming Huang1,3, [^1] Co-first authors with equal contribution., Yibo Yan1,2, [^1] Co-first authors with equal contribution., Jiahao Huo1, Xin Zou1, Xinfeng Li4, Kun Wang4, Xuming Hu1,2, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University {hmhuang04, yanyibo70}@gmail.com, {xuminghu}@hkust-gz.edu.cn
@@ 2.2 Knowledge Circuit Analysis
In the context of mechanistic interpretability Rai et al. (2024); Huo et al. (2024); Huang et al. (2024a), computations in Transformer-based language models are viewed as a connected directed acyclic graph encompassing components such as MLPs and attention layers Syed et al. (2023); Conmy et al. (2023); Huang et al. (2024b). A circuit refers to a sparse computational subgraph that significantly influences the model’s behavior on a specific task Olah et al. (2020); Elhage et al. (2021); Wang et al. (2022). Building on this, Yao et al. (2024) introduce the concept of knowledge circuits, hypothesizing that cooperation among components reveals implicit knowledge in LLMs. Further, Ou et al. (2025); Zucchet et al. (2025); Hakimi et al. (2025) explore how such circuits evolve during continual pre-training, providing insights into knowledge acquisition. To enable effective knowledge editing, CaKE Yao et al. (2024) proposes a Circuit-aware Knowledge Editing method that guides models to activate modified knowledge and form new reasoning pathways. In this paper, we analyze the phenomenon of knowledge overshadowing through the lens of knowledge circuits, contributing new perspectives to LLM hallucination detection.
@@
Figure 2: A group of imbalanced natural language data of knowledge overshadowing. The $X_{bg}$, $X_{dom}$ and $X_{sub}$ are background, dominant and subordinate knowledge respectively. The $P_{dom}$ and $P_{sub}$ are prompt with dominant and subordinate knowledge respectively, their corresponding answers are $Y_{dom}$ and $Y_{sub}$.
@@ 3.1 Knowledge Overshadowing
As shown in Figure 2, let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$. The core idea is that a strong learned pattern $X_{dom}\leftrightarrow X_{bg}$ can "over-generalize" where the model primarily associates $X_{dom}$ with $X_{bg}$. Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$, it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
@@ 3.1.1 Knowledge Overshadowing Occurrence
When the input prompt is $P_{sub}$, but the model wrongly generates $Y_{dom}$, knowledge overshadowing occurs for this query-response instance, resulting in the $(P_{sub},Y_{dom})$. In the case shown in Figure 2, if overshadowing occurs, the model will incorrectly answer "Beijing" to the input prompt "Where is the most important economic center of China?".
@@ 3.1.2 Quantitative Indicators
Let $N_{dom}$ be the number of instances of the dominant knowledge prompt $P_{dom}$ and $N_{sub}$ be the number of instances of the subordinate knowledge prompt $P_{sub}$ in a group of data, respectively. In the case shown in Figure 2, $N_{dom}$ and $N_{sub}$ are 2 and 1, respectively. The entire dataset might contain multiple different groups of imbalanced data.
@@ 3.1.3 Overshadowing Influence Factors
Knowledge Popularity (P) is the fundamental cause of the knowledge overshadowing phenomenon and serves as its primary influencing factor. P is defined as the ratio of $N_{dom}$ to $N_{sub}$ in a group,
$P=\frac{N_{dom}}{N_{sub}}.$ (4)
In the case of Figure 2, $P=2$. A higher value of P signifies a more severe data imbalance.
@@
In addition to the factors mentioned in Zhang et al. (2025a, 2024b), aiming to analyze the dynamic evolution of knowledge overshadowing during the training process, we extend our consideration to total number of tokens, the Dataset Size (D) in the training set. Furthermore, the average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch during training is defined as
$\mathcal{LP}=\frac{loss(P_{sub})}{total\;loss},$ (5)
which measures the allocation of the model’s efforts for the optimization of subordinate knowledge and recovery from knowledge overshadowing.
@@ 3.2.2 Knowledge Circuit Construction
We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs. The node set $V$ encompasses input embeddings, attention heads, MLP layers, and output logits. The edge set $E$ represents the information flow between these components. Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of given input prompt, particularly in knowledge overshadowing, is $\{X_{dom},X_{sub}\}$, the difference between $P_{dom}$ and $P_{sub}$. The adapted construction method is similar to Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al. (2024), which involves:
Paired inputs. For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrast mentioned below.
Activation difference calculation. After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activations of the parent node for the clean input $P_{sub}$ and the corrupt input $P_{dom}$, respectively.
Edge score calculation. The importance $S(e)$ of an edge $e=(v_{p},v_{c})$ is scored based on how patching $v_{p}$’s activation (using $\Delta A(v_{p})$) influences a metric $\mathcal{M}$, which assesses the model’s ability to correctly output $Y_{sub}$ rather than $Y_{dom}$ for $P_{sub}$ inputs. Using the Integrated Gradients (IG), $S(e)$ is approximated as:
@@
Circuit in Knowledge Overshadowing. In the context of knowledge overshadowing, the metric $\mathcal{M}$ is defined as the difference between the output logits of $Y_{sub}$ and $Y_{dom}$. A positive value for $\mathcal{M}$ indicates that the model correctly predicts $Y_{sub}$. Therefore, we can represent the circuit where overshadowing occurs as $C_{overshadowed}=\{V,E\}$ with $\mathcal{M}<0$ and $S(E|P_{sub},Y_{dom})\geq\tau$. Furthermore, the competition between the dominant and subordinate knowledge pairs during circuit construction is quantified by the activation values of different input $P_{dom}$ and $P_{sub}$, $\Delta A(v_{p})=A_{P_{sub}}(v_{p})-A_{P_{dom}}(v_{p})$.
@@ 3.2.3 Circuit-based Analysis
Node Attention Analysis. We identify high-attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
Circuit Structure Analysis. We then trace the information flow of these high-attention heads by identifying their parent and child nodes to understand their structural role. Nodes consistently retained in circuits built with different thresholds $\tau$ are also analyzed as key components.
@@ 3.2.4 Circuit-based Overshadowing Recovery
where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$. We use a two-stage optimization method to find the optimal number of edges. First, we build the circuit by adding edges at regular intervals, revealing that the circuit’s performance exhibits a multi-modal relationship with the edge count. Subsequently, we apply the golden-section search algorithm within the optimal range to determine the $\tau_{opt}$ for the $C_{opt}$. This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
@@
Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, are masked. $Y_{dom}$ is identified as $y_{i}$ that has the highest average rank across all $P^{\prime}_{sub}$. For circuit construction, $X_{dom}$ within $P_{dom}$ is replaced by a generic placeholder like "something", as shown in Figure 6. Combining this streamlined approach enables broader application of our knowledge circuit-based overshadowing recovery. See more details in Appendix B.2
@@ 4.1 Experiment Setup
Figure 4: The co-evolution between $\mathcal{RO}$ & $\mathcal{LP}$ in different M. Early high overall loss and low $\mathcal{LP}$ leads to intensive optimization of $P_{dom}$ and $\mathcal{RO}$ rises up to near 100%. As training progresses, subordinate knowledge loss proportion ($\mathcal{LP}$) rises, shifting optimization focus to $P_{sub}$ errors, initiating $\mathcal{RO}$’s recovery phase, validated across models.
@@ 4.1.1 Dataset
fine-tuning Dataset. We construct a fine-tuning dataset to evaluate circuit-base overshadowing recovery method. Utilizing virtual knowledge, we preserve the natural language semantics while avoiding prior knowledge editing that could interfere with the natural occurrence of overshadowing. For this dataset, we set P = 5 and D = 1M. The Appendix D shows the dynamics analysis based on fine-tuning dataset.
@@ 4.1.3 Training
Figure 5: The circuit-based analysis of proposed PhantomCircuit. (a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch. (b) focuses on the onset and recovery phases and shows that the appear and disappear of high-attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$. (C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers. The main structure of circuit (with 400 edges totally) and attention mode show that high-attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank. Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
@@ 4.1.4 Evaluation
To measure $\mathcal{RO}$, we randomly sample 500 $P_{dom}$ and 500 $P_{sub}$ prompts for evaluation after each training epoch. The $\mathcal{LP}$ is recorded within each epoch. The results are shown in Figures 3 and 4.
@@ 4.2 Main Result
As shown in Figures 3(a) and 3(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase. This is attributed to more prominent dominant patterns $X_{dom}\leftrightarrow X_{bg}$ being learned and generalized rapidly, even within the first epoch. The duration phase also decreases with higher P, because a larger P and a fixed D implies fewer groups of knowledge pairs, leading to less diversity of $P_{sub}$, allowing the model to learn all overshadowed $P_{sub}$ and recover from overshadowing more quickly.
Figures 3(b) and 3(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing. This is due to the stronger generalization capabilities of larger models, leading them to quickly learn and overgeneralize dominant patterns. However, larger models also show a shortened duration phase and a rapid decline in $\mathcal{RO}$ during recovery, suggesting enhanced capacity to differentiate $\{X_{dom},X_{sub}\}$, thus recovering faster despite earlier overshadowing.
Figure 6: Overshadowing recovery via optimized circuit. In the fine-tuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$. First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ . Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
In Figures 3(c) and 3(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner. This is because more data per epoch provides more iterations and exposure to the dominant pattern, accelerating its generalization. Conversely, the recovery phase is prolonged with larger datasets. The increased diversity of $P_{sub}$ in larger datasets requires more epochs for the model to learn all instances and recover.
@@
The occurrence of overshadowing in pre-trained LLMs can be understood with dynamics analysis results above. Initially, large M and D promote the rapid generalization of $X_{dom}\leftrightarrow X_{bg}$ and overshadowing onset. The subsequent inadequate optimization of the $P_{sub}$ is exacerbated by the sheer scale and diversity of training data, which brings prolonged overshadowing effect and exceeded sharp recovery effect from large M (e.g., trillion tokens for training Llama-2-7B).
@@
The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs influences the extent of knowledge overshadowing. Figure 5 (a) shows the variation of circuit’s average attention scores on $\{X_{dom},X_{sub}\}$ throughout the training phase. The higher average attention alleviates overshadowing, while lower attention exacerbates it.
@@
Figure 5 (b) shows the circuit dynamics across the onset and recovery phases of knowledge overshadowing. These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high-attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit. The threshold for high-attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset. Conversely, when knowledge overshadowing intensifies, a subset of these critical attention heads tends to disappear from the circuit.
Table 1: Ablation Study for high-attention Heads on fine-tuning Dataset. The model’s performance degrades with the ablation of high-attention heads, which validates the circuit’s faithfulness.
Model Proportion of Ablated High Attn. Heads 10% 20% 50%
gpt2-medium Performance Degradation 1.03 1.43 1.67
Attn. Score on $X_{sub}$ Degradation 0.044 0.086 0.141
pythia-410m Performance Degradation 0.92 1.30 1.42
Attn. Score on $X_{sub}$ Degradation 0.039 0.074 0.126
@@
Furthermore, by focusing on the circuit’s internal mechanisms and structure within a specific epoch, as shown in Figure 5 (c), we leverage the circuit-based analysis of our proposed PhantomCircuit. First, according to the logits and ranks of $Y_{dom}$ and $Y_{sub}$ across layers, the 5th layer is a juncture where the rank of $Y_{sub}$ exceeds $Y_{dom}$. Subsequently, by focusing on the circuit structure, we identify the internal mechanisms driving this juncture. A high-attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5. The a5.h1 also appears to inherit its high-attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal. The attention map on the right confirms its high focus on $\{X_{dom},X_{sub}\}$.
@@
Knowledge circuit-guided optimization represents a promising strategy to mitigate the knowledge overshadowing effect. Figure 6 illustrates our findings. We first fine-tune the model on fine-tuning dataset. During the recovery phase, we randomly choose some $P_{sub}$. Feeding $P_{sub}$ to the original full model (M = 410M), the prediction is still $Y_{dom}$ instead of expected $Y_{sub}$, and shows a significant gap in probability as well.
@@
We then detect the $Y_{sub},Y_{dom}$ and $X_{sub}$ for chosen $P_{sub}$ automatically, replace the $X_{sub}$ of placeholder token "something". With these components, we obtain the $P_{dom}$ as corrupt input and $P_{sub}$ as clean input to construct a knowledge circuit. The optimized circuit graph shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information. Retained high-attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$. Some low attention heads also remain, implying that even in the circuit, processing background knowledge $X_{bg}$ and linking it to the distinctive elements $X_{sub}$ is crucial for correct inference. The performance of the optimized circuit $C_{opt}$ is then evaluated by feeding it the clean input $P_{sub}$, while $P_{dom}$ serves as the baseline for contrast. Finally, the circuit successfully produces $Y_{sub}$, demonstrating the elimination of the overshadowing effect.
4.3 Circuit Faithfulness Study
Based on the work of Marks et al. (2024); Hanna et al. (2024), the EAP-IG method builds circuits with high faithfulness and efficiency, ensuring our findings based on circuit analysis are reliable. To verify this key property in the context of knowledge overshadowing, we conduct an ablation study on high-attention heads.
We first construct optimal circuits for pythia-410m and gpt2-medium models trained on the fine-tuning dataset. Subsequently, we removed the top 10%, 20%, and 50% of attention heads by score and observed the impact on performance metric $\mathcal{M}$ and attention scores, with results shown in Table 1.
The results show that ablating just 10% of the high-attention heads leads to a significant drop in both the attention scores on $X_{dom},X_{sub}$} and the circuit performance metric $\mathcal{M}$. This performance degradation is substantial enough to alter the output in many cases, preventing the model from accurately predicting $Y_{sub}$. In contrast, at 50% ablation, the drop in $\mathcal{M}$ becomes less pronounced because the model’s ability to predict both the $Y_{sub}$ and $Y_{dom}$ severely declines. These findings validate the critical impact of the nodes in optimal circuits, especially high-attention heads, on model performance and confirm the circuit’s faithfulness.